A and B are partners in a firm sharing profits in the ratio of 2 : 1. C joins the firm.
A surrenders 1/4th of his share and B 1/5th of his share in favour of C. Find the new profit sharing ratio.
1
Expert's answer
2019-08-23T10:58:54-0400

New Ratio = Old Ratio - Sartificing Ratio

A`s = 2/3-1/4=25/60

B`s = 1/3-1/5=8/60

C`s = 60/60-25/60-8/60 = 27/60

New Profit Ratio = 25:8:27